Let the population be given by the numbers. Take all random samples of size 3 without replacement and obtain the sampling distribution of the sample mean


1
Expert's answer
2021-04-29T10:58:43-0400

Let the population consists of numbers "15,9,24,6" and "18" .

Sampling distribution of size 3 can be given as




Sampling distribution of Sample mean,

"P(\bar X = 16) = \dfrac{3}{10}"


"P(\bar X = 15) = \dfrac{1}{10}"


"P(\bar X = 19) = \dfrac{3}{10}"


"P(\bar X = 13) = \dfrac{1}{10}"


"P(\bar X = 11) = \dfrac{1}{10}"


"P(\bar X = 14) = \dfrac{1}{10}"